### v1

# SegmentDreamer: Towards High-fidelity Text-to-3D Synthesis with Segmented Consistency Trajectory Distillation

Abstract
Recent advancements in text-to-3D generation improve the visual quality of Score Distillation Sampling (SDS) and its variants by directly connecting Consistency Distillation (CD) to score distillation. However, due to the imbalance between self-consistency and cross-consistency, these CD-based methods inherently suffer from improper conditional guidance, leading to sub-optimal generation results. To address this issue, we present SegmentDreamer, a novel framework designed to fully unleash the potential of consistency models for high-fidelity text-to-3D generation. Specifically, we reformulate SDS through the proposed Segmented Consistency Trajectory Distillation (SCTD), effectively mitigating the imbalance issues by explicitly defining the relationship between self- and cross-consistency. Moreover, SCTD partitions the Probability Flow Ordinary Differential Equation (PF-ODE) trajectory into multiple sub-trajectories and ensures consistency within each segment, which can theoretically provide a significantly tighter upper bound on distillation error. Additionally, we propose a distillation pipeline for a more swift and stable generation. Extensive experiments demonstrate that our SegmentDreamer outperforms state-of-the-art methods in visual quality, enabling high-fidelity 3D asset creation through 3D Gaussian Splatting (3DGS).
Figure 1: Examples of 3D assets generated by SegmentDreamer. Our framework addresses the improper conditional guidance issues of recent Consistency Distillation (CD)-based methods and theoretically provides a significantly tighter upper bound on distillation error, enabling high-fidelity text-to-3D generation within a reasonable timeframe ($\sim$32 minutes with classifier-free guidance [9] and $\sim$38 minutes combined with Perp-Neg [1]) on a single A100 GPU through 3D Gaussian Splatting (3DGS). Project Page: https://zjhJOJO.github.io
1 Introduction
Text-to-3D generation aims to create 3D assets that accurately align with text descriptions. It plays a vital role in various domains such as virtual reality, 3D gaming, and film production. One compelling solution is to directly “distill” 3D assets from 2D text-to-image Diffusion Models (DM) [34] through Score Distillation Sampling (SDS) [31, 46], which leverages pre-trained text-to-image DMs as the 3D priors to guide the optimization of differentiable 3D representations, such as Neural Radiance Fields (NeRF) [29] and 3D Gaussian Splatting (3DGS) [17]. Subsequent research focus on improving the performance of SDS in diversity [47, 11, 43], visual quality [3, 24, 47, 51, 58, 48, 22, 25, 53, 57, 16], and multi-view consistency [5, 10, 37, 35, 12].
Figure 2: Visual example of existing CD-based methods: CDS [48], GCS [22], and our SegmentDreamer with a Classifier-Free Guidance (CFG) scale of 7.5. As shown, improper conditional guidance leads to suboptimal results. (a) Ineffective guidance produces semantically inconsistent details, while (b) excessive guidance results in overexposure and artifacts. Although (c) Brightness-Equalized Generation (BEG) [22] can mitigate overexposure, artifacts persist. In contrast, our SegmentDreamer enhances conditional generation, producing high-fidelity 3D assets.
Recent methods [48, 22] have improved SDS and its variants by incorporating Consistency Distillation (CD) [40, 18] into their distillation loss. However, due to an imbalance between self-consistency and cross-consistency in their losses, these methods inherently suffer from improper conditional guidance when combined with Classifier-Free Guidance (CFG) for generation, ultimately leading to suboptimal outcomes (see Figure 2, a detailed explanation is provided in Section 4.1). Additionally, the distillation error upper bound of these methods remains large, limiting their generation quality.
To address the aforementioned limitations, we present a novel framework – SegmentDreamer, which aims to fully unleash the potential of consistency models for high-fidelity text-to-3D generation. Specifically, by analyzing existing CD-based methods, we draw inspiration from segmented consistency models [33, 44] to reformulate SDS through the proposed Segmented Consistency Trajectory Distillation (SCTD), which effectively mitigates the imbalance between self-consistency and cross-consistency by explicitly defining their relationship. Moreover, unlike CD-based methods that directly enforce consistency across the entire Probability Flow Ordinary Differential Equation (PF-ODE) trajectory, SCTD partitions the PF-ODE trajectory into multiple sub-trajectories, ensuring consistency within each segment. This approach theoretically provides a significantly tighter upper bound on distillation error. In addition, we propose a distillation pipeline for a more swift and stable generation. We also introduce 3DGS as our 3D representation. Extensive experimental results demonstrate that our SegmentDreamer outperforms state-of-the-art methods in visual quality. Moreover, the proposed SCTD can be seamlessly applied to various 3D generative tasks (e.g., 3D avatar generation, and 3D portrait generation), facilitating the creation of high-fidelity 3D assets.
The main contributions are summarized as follows:
  • We present SegmentDreamer, a novel framework for high-fidelity text-to-3D generation, which addresses the imbalance issue of existing CD-based methods.
  • We reformulate SDS through the proposed SCTD, effectively mitigating the imbalance between self- and cross-consistency. SCTD also theoretically provides a significantly tighter upper bound on distillation error. We further propose a distillation pipeline for more swift and stable optimization.
  • Experimental results demonstrate that our SegmentDreamer outperforms state-of-the-art methods in visual quality, enabling high-fidelity 3D asset generation through 3DGS.
2 Related Work
2.1 Text-to-3D Generation
Creating a 3D object from a given text description is a long-standing challenge in the fields of computer vision and graphics. To achieve this goal, a series of end-to-end text-to-3D pipelines have been proposed [8, 7, 30, 13, 21, 49], but these methods require large-scale labeled 3D datasets for training. To reduce the reliance degree of 3D training data, a compelling solution is to directly “distill” 3D assets from 2D text-to-image Diffusion Models (DM) [34]. A groundbreaking contribution in this domain is DreamFusion [31], which introduced Score Distillation Sampling (SDS) loss to optimize parametric 3D representations. Follow-up works focus on refining and extending SDS in various aspects. For example, several studies [47, 11, 43] improve SDS in diversity through prompting modifications, while [3, 24, 47, 51, 58, 48, 22, 25, 53, 57, 16] enhance the fidelity of the generated 3D objects. Other advancements aim to achieve multi-view consistency [5, 10, 37, 35, 12] and faster generation [42, 52]. Recently, Consistency Distillation (CD) [40, 18] has received widespread attentions as it enables high-quality generation through minimal sampling steps. Inspired by CD, Wu et al. [48] and Li et al. [22] incorporate CD into their distillation loss, which addresses some limitations brought by SDS. However, these CD-based methods inherently suffer from improper conditional guidance, ultimately leading to sub-optimal results. In this work, we aim to unleash the potential of CD to generate high-fidelity 3D assets, pushing the boundaries of text-to-3D synthesis.
2.2 Differentiable 3D Representations
Neural Radiance Field (NeRF) [29] aims to construct implicit 3D representations using a multi-layer perceptron (MLP), leveraging standard volumetric rendering [14] and alpha compositing techniques [32]. Subsequent research has explored adapting NeRF to various applications, Subsequent research has explored adapting NeRF to various applications, including sparse-view reconstruction [45], anti-aliasing [2], and medical image super-resolution [4].
3D Gaussian Splatting (3DGS) [17] has emerged as a leading approach for 3D representation, offering rapid optimization and high-speed rendering. Unlike NeRF, 3DGS explicitly models scenes using a set of 3D Gaussians and renders views through splatting [20]. Its versatility has enabled applications across various domains, including avatar generation [36, 55], single-view reconstruction [41], anti-aliasing [54], digital watermarking [6], and SLAM [28, 50].
3 Preliminaries
Diffusion Model (DM) [39, 38, 15] consists of a diffusion process and a denoising process. The diffusion process can be modeled as the solution $\hat{\text{o}}$ stochastic differential equation (SDE) with the transition kernel $\mathbb{P}(\mathbf{z}_{t}|\mathbf{z}_{0}):=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}\mathbf{z}_{0},\sigma_{t}\mathbf{I})$. By reversing the diffusion process, $\mathbf{z}_{0}$ can be estimated by solving a Probability Flow Ordinary Differentiable Equation (PF-ODE), which is defined by $\mathrm{d}\mathbf{z}_{t}=[f(t)\mathbf{z}_{t}-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{z}_{t}}\mathrm{log}\mathbb{P}(\mathbf{z}_{t})]\mathrm{d}t$. A neural network ${\boldsymbol{\epsilon_{\phi}}}(\mathbf{z},t)$ can be trained to approximate the scaled score function $-\sigma_{t}\nabla_{\mathbf{z}_{t}}\mathrm{log}\mathbb{P}(\mathbf{z}_{t})$, offering an empirical formulation of the PF-ODE: $\mathrm{d}\mathbf{z}_{t}=[f(t)\mathbf{z}_{t}+\frac{g(t)^{2}}{2\sigma_{t}}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t)]\mathrm{d}t.$
Consistency Model [40, 27] aims to achieve few-step inference by distilling a pre-trained DM $\boldsymbol{\phi}$ into a student model $\boldsymbol{\phi}^{\prime}$. Given a PF-ODE trajectory $\{\mathbf{z}_{t}\}_{t\in[0,T]}$ and a condition $\mathbf{y}$, a consistency model $\boldsymbol{f}_{\boldsymbol{\phi}^{\prime}}$ builds the mapping $\boldsymbol{f}_{\boldsymbol{\phi}^{\prime}}(\mathbf{z}_{t},t,\mathbf{y})=\mathbf{z}_{0}$ between any points along this trajectory and the endpoint $\mathbf{z}_{0}$ by minimizing the following Consistency Distillation (CD) loss:
$\mathcal{L}_{\text{CD}}=\mathbb{E}[\lambda(t)||\boldsymbol{f}_{\boldsymbol{\phi}^{\prime}}(\mathbf{z}_{t},t),\boldsymbol{f_{\phi_{s}^{-}}}(\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s},s)||_{2}^{2}],t>s,$ (1)
where $\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ for $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\boldsymbol{\phi_{s}^{-}}$ is updated with the exponential moving average [40, 27]. $\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}$ can be obtained by:
$\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}$ $=\boldsymbol{\Phi}(\mathbf{z}_{t},t,s,\mathbf{y})$ (2)
$=\mathbf{z}_{t}+\int_{t}^{s}[f(u)\mathbf{z}_{u}+\frac{g(u)^{2}}{2\sigma_{u}}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{u},u,\mathbf{y})]\mathrm{d}u.$
The consistency function $\boldsymbol{f_{\phi_{s}}}$ can be parameterized in various ways [40, 27], and the solver $\boldsymbol{\Phi}(\cdot,\cdot,\cdot,\cdot)$ is used to calculate the point at $s$, which can be implemented using DDIM [38] or DPM-solver [26].
Score Distillation Sampling (SDS) Loss. Given a 3D representation $\boldsymbol{\theta}$ that can render a view image $\mathbf{z}_{0}:=\boldsymbol{g}(\mathbf{c},\boldsymbol{\theta})$ with camera pose $\mathbf{c}$, SDS distills a 2D DM $\boldsymbol{\phi}$ to generate a 3D object. by minimizing the following SDS loss w.r.t. $\boldsymbol{\theta}$:
$\mathcal{L}_{\text{SDS}}:=\mathbb{E}_{t,\boldsymbol{\epsilon}}[\omega(t)||\underbrace{\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}}_{\text{noise residual}}||_{2}^{2}],$ (3)
where $\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$. Poole et al. [31] empirically omit the U-Net Jacobian term $\frac{\partial\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})}{\mathbf{z}_{t}}$ in gradient calculation for speed-up, leading to low-fidelity results.
4 Our Method
In this section, we first review the mechanism of existing Consistency Distillation (CD)-based methods [48, 22] and reveal that they inherently suffer from improper conditional guidance when combined with Classifier-free Guidance (CFG). Subsequently, we address these limitations by reformulating SDS through the lens of the proposed Segmented Consistency Trajectory Distillation (SCTD). Moreover, building on this insight, we propose the efficient SCTD sampling for high-fidelity text-to-3D generation. We further tailor a distillation pipeline for the proposed SCTD sampling to achieve more swift and stable optimization. Finally, we provide a theoretical proof for our distillation error upper bound. The overall framework of our SegmentDreamer is depicted in Fig. 3, with the algorithm detailed in Algorithm 1.
4.1 Reviewing CD-based Losses
The consistency function w.r.t. 3D representation $\boldsymbol{\theta}$ is parameterized by $\boldsymbol{F_{\theta}}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\mathbf{z}_{t}-\sigma_{t}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})}{\alpha_{t}}$ [48, 22]. Before proceeding, we clarify two key concepts: self-consistency and cross-consistency. The former ensures that all points on the same ODE trajectory consistently map to their original points, as in consistency models [27, 40, 18]. The latter enforces an alignment between unconditional and conditional ODE trajectories [22].
Consistency Distillation Sampling (CDS) loss [48] is defined as
$\mathcal{L}_{\text{CDS}}(\boldsymbol{\theta})=\mathbb{E}_{t,s}[c(t)||\boldsymbol{F_{\theta}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{F_{\theta}}(\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s},s,\mathbf{y})||_{2}^{2}],$ (4)
where $s<t$, $c(t)=\omega(t)(\frac{\alpha_{t}}{\sigma_{t}})^{2}$, and $\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}$ is estimated using single-step DDIM sampling. Eq. 4 reveals that CDS prioritizes self-consistency while omitting cross-consistency in its loss function, which can be interpreted as the absence of the “classifier score”. As reported in CSD [53], this missing component, crucial for conditional generation, hinders CDS from producing semantically consistent content (see Fig. 5 for details). Moreover, the reliance on $\boldsymbol{F}_{\theta}$ may introduce a significant distillation error upper bound $\mathcal{O}(\Delta_{t})T$, where $\Delta_{t}=\max{|t-s|}$. This issue is particularly pronounced in text-to-3D settings due to the typically large $|t-s|$.
Guided Consistency Sampling (GCS) [22] leverages the DPM-Solver-1 [26] as its consistency function:
$\boldsymbol{G_{\theta}}(\mathbf{z}_{t},t,s,\mathbf{y})=\frac{\alpha_{s}}{\alpha_{t}}\mathbf{z}_{t}-\alpha_{s}\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})\int_{\lambda_{t}}^{\lambda_{s}}e^{-\lambda}\mathrm{d}\lambda,$ (5)
where $\lambda_{t}=\mathrm{ln}\frac{\alpha_{t}}{\sigma_{t}}$. Based on Eq. 5, GCS proposes a more compact consistency (CC) loss $\mathcal{L}_{\text{CC}}$ defined as
$\mathcal{L}_{\text{CC}}(\boldsymbol{\theta})\!=\!\mathbb{E}_{t,s,e}[\left\|\boldsymbol{G_{\theta}}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,e,\emptyset)\!-\!\boldsymbol{G_{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,e,\emptyset)\right\|_{2}^{2}],$ (6)
where $e$$<$$s$$<$$t$. Moreover, GCS employs a conditional guidance loss $\mathcal{L}_{\text{CG}}$ and a pixel-wise constraint $\mathcal{L}_{\text{CP}}$ to enforce cross-consistency (see Supp. A for details).
However, GCS loss suffers from improper conditional guidance due to two key issues: (1) $\boldsymbol{G_{\theta}}$ used to enforce cross-consistency is inherently flawed due to the absence of a target time-step in $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})$, as detailed in App. B. (2) GCS employs $\boldsymbol{F}_{\boldsymbol{\theta}}$, implemented via one-step DDIM, to enforce cross-consistency throughout the ODE trajectory in both latent and pixel spaces. However, as reported in its paper, it suffers from unstable generation due to the excessive conditional guidance, leading to overexposure and artifacts. Furthermore, this approach results in a less stringent upper bound on the distillation error, expressed as $\mathcal{O}(\Delta_{t})(T-e)$, which constrains its generation quality.
Summary: The improper conditional guidance in CDS and GCS arises mainly from an imbalance between self-consistency and cross-consistency. Moreover, the reliance on $\boldsymbol{F_{\theta}}$ to enforce consistency across the entire trajectory further contributes to this issue.
4.2 Reframing SDS through SCTD
We address the previous limitations by reframing SDS through the lens of SCTD. Specifically, we first define our consistency function as follows: as existing segmented consistency models [44, 33] do, we partition the entire time-step range $[0,T]$ into $N_{s}$ sub-intervals (i.e., $N_{s}+1$ edge points), represented by $\{[s_{m},s_{m+1})|m\in\{0,\cdots,N_{s}-1\},s_{0}=0,s_{N_{s}}=T\}$ (the segmentation strategy will be discussed Sec. 4.3). We define our consistency function $\boldsymbol{G_{\theta}}$ using Eq. 5, aiming to enforce consistency within each sub-trajectory. i.e., $\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,s_{m},\mathbf{y})=\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{s},s,s_{m},\mathbf{y})$ for $s_{m}\leq s<t\leq s_{m+1}$. For brevity, we abbreviate $\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,s_{m},\mathbf{y})$ as $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\mathbf{z}_{t},t,\mathbf{y})$. Notably, our consistency function $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}$ avoids the issues inherent in GCS, as analyzed in Sec. 4.1 and Supp. B, since at each training stage, the target time step $s_{m}$ is always uniquely determined.
When CFG [9] with scale $\omega$ is used, the noise residual term in Eq. 3 can be transformed into
$\hat{\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}$ $=\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}$ (7)
$+\omega(\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\emptyset)).$
Based on Eq. 7, $\mathcal{L}_{\text{SDS}}$ can be expressed in the form of SCTD:
$\mathcal{L}_{\text{SDS}}(\boldsymbol{\theta})=\mathbb{E}_{t,s}[b(t)||\underbrace{\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\mathbf{y})}_{\text{self-consistency constraint}}+$ (8)
$\omega\underbrace{(\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\emptyset)\!-\!\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\mathbf{y}))}_{\text{cross-consistency constraint}}\!+\!\underbrace{\mathbf{z}_{s_{m}}-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})}_{\text{generative prior}}||_{2}^{2}].$
where $b(t)=\frac{\omega(t)}{(\alpha_{s_{m}}\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda)^{2}}$ and $\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\mathbf{z}_{t},t,s,\mathbf{y})$. The derivation is given in App. C. The first term on the r.h.s. of Eq. 8 enforces self-consistency across the conditional denoising sub-trajectory, referred to as the “self-consistency constraint”. The second term enforces cross-consistency between the unconditional and conditional denoising sub-trajectories, termed the “cross-consistency constraint”. The third term adheres to the naming conventions in CSD [53].
As observed, Eq. 8 explicitly defines the relationship between the self-consistency constraint and the cross-consistency constraint, effectively addressing the imbalance problem present in previous CD-based methods [48, 22]. Furthermore, as demonstrated in Sec. 4.3, SCTD provides a tighter upper bound on distillation error, enhancing the upper limit of our algorithm.
Figure 3: An overview of SegmentDreamer: We begin by initializing a 3D representation $\boldsymbol{\theta}$ using a 3D generator, such as Point-E [30]. In each iteration, we randomly render a batch of camera views $\mathbf{z}_{0}$ from $\boldsymbol{\theta}$ and diffuse them into $\mathbf{z}_{s_{m}}$ with fixed noise $\boldsymbol{\epsilon}^{*}$. Next, we transform $\mathbf{z}_{s_{m}}$ into $\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t}$ using either one-step or two-step unconditional deterministic sampling. During the denoising process, we first estimate $\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}$ through one-step conditional deterministic sampling from $\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t}$. Subsequently, we compute two consistency functions and utilize them to derive the loss $\mathcal{L}_{\text{SCTD}}$, which is ultimately employed to optimize $\boldsymbol{\theta}$.
4.3 SCTD Sampling
Sampling Design. A natural question arises: according to Eq. 8, does minimizing $\mathcal{L}_{SDS}$ necessarily enforce both self- and cross-consistency? The answer is no. This can be easily understood by noting that one cannot necessarily achieve $x_{1}=x_{2}$ and $y_{1}=y_{2}$ when $||x_{1}-x_{2}+\omega(y_{1}-y_{2})||_{2}^{2}=0$. While Eq. 8 provides valuable insights, applying it to text-to-3D tasks remains challenging. To address this, we introduce the following modifications. First, we observe that the condition $\mathbf{y}$ appears in all three loss terms in Eq. 8, which increases computational cost particularly when using Perp-Neg [1] or ControlNet [56]. To mitigate this, we apply an equivalent transformation to it:
$\mathcal{L}_{\text{SDS}}(\boldsymbol{\theta})=\mathbb{E}_{t,s}[b(t)||\underbrace{\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\emptyset)-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\emptyset)}_{\text{self-consistency constraint}}+$ (9)
$(\omega\!+\!1)\!\underbrace{(\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\emptyset)\!-\!\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\mathbf{y}))}_{\text{cross-consistency constraint}}\!+\!\underbrace{\mathbf{z}_{s_{m}}\!-\!\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\emptyset)}_{\text{generative prior}}\!||_{2}^{2}].$
The derivation is given in Supp. D.
We follow the suggestion of CSD [53], omitting the generative prior from Eq. 9. To enforce the self- and cross-consistency, we consider a more strict constraint on Eq. 8, defining our SCTD sampling as follows:
$\mathcal{L}_{\text{SCTD}}(\boldsymbol{\theta})=\mathbb{E}_{t,s}[b(t)[||\underbrace{sg(\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\emptyset))-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\emptyset)}_{\text{self-consistency constraint}}||_{2}^{2}$ (10)
$+||(\omega+1)\underbrace{(\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\emptyset)-sg(\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})))}_{\text{cross-consistency constraint}}||_{2}^{2}]].$
where $\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}}$ is obtained through multi-step deterministic sampling from $\mathbf{z}_{0}$, which is discussed in Sec. 4.4. $\mathbf{z}_{s_{m}}=\alpha_{s_{m}}\mathbf{z}_{0}+\sigma_{s_{m}}\boldsymbol{\epsilon}^{*}$ with $\boldsymbol{\epsilon}^{*}$ fixed for the entire training. $\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,s,\mathbf{y})$ where $\boldsymbol{\Phi}$ is implemented by a first-order ODE solver in this work. $sg(\cdot)$ represents stopping gradient backpropagation.
Algorithm 1 Text-to-3D Generation via SCTD Loss
1:Initialization: threshold $t_{\tau}$, segmentation number $N_{s}$, and target prompt $\mathbf{y}$
2:Initialize $\boldsymbol{\theta}$ with a pre-trained 3D generator
3:while $\boldsymbol{\theta}$ is not converged do
4:     Sample $\mathbf{z}_{0}=\boldsymbol{g}(\mathbf{c},\boldsymbol{\theta})$ and $t\sim\mathcal{U}(0,T)$
5:     Obtain $s_{m}\!=\!max\{s_{i}|s_{i}\leq t,i=0,\cdots,N_{s}-1\}$
6:     Obtain $\mathbf{z}_{s_{m}}=\alpha_{s_{m}}\mathbf{z}_{0}+\sigma_{s_{m}}\boldsymbol{\epsilon}^{*}$
7:     Sample $s\sim\mathcal{U}(s_{m},t)$
8:     Predict $\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}}$ with Eq. 11 and $\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}\!=\!\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},t,s,\mathbf{y})$
9:     Obtain $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\emptyset)$,$\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$ with Eq. 5
10:     Obtain $\nabla_{\theta}\mathcal{L}_{\text{SCTD}}$ with Eq. 10 and update $\boldsymbol{\theta}$.
11:end while
Trajectory Segmentation Strategy. We propose two heuristic trajectory segmentation strategies: equal segmentation and monotonically increasing segmentation. In the first strategy, the entire time-step range $[0,T]$ is uniformly divided into $N_{s}$ segments, each with an equal duration, i.e., $s_{m+1}-s_{m}=\frac{T}{N_{s}}$. In the second strategy, we adopt a monotonically increasing interval segmentation method, where the $N_{s}$ intervals satisfy the condition: $s_{m+1}-s_{m}=t_{\tau}+\frac{2m(T-N_{s}t_{\tau})}{N_{s}(N_{s}-1)},m\in\{0,\cdots,N_{s}-1\}$. A comparison of these two strategies is provided in Sec. 6.
4.4 Swift and Stable Optimization Pipeline
In previous work [23, 48, 22], $\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}}$ is derived from deterministic sampling with a fixed number of steps, lacking flexibility. In this paper, we adopt the following dynamic adjustment strategy: .
$\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}}=\begin{cases}\boldsymbol{\Phi}(\boldsymbol{\Phi}(\mathbf{z}_{s_{m}},s_{m},s,\emptyset),s_{m},t,\emptyset),&\text{if }t>t_{\tau}\\ \boldsymbol{\Phi}(\mathbf{z}_{s_{m}},s_{m},t,\emptyset),&\text{if }t\leq t_{\tau}\end{cases},$ (11)
This strategy enable us to make a trade-off between generation quality and optimization speed, as evidenced in Sec. 6.
Consistency function Approximation. For faster optimization, we propose approximating $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\emptyset)$ as $\mathbf{z}_{s_{m}}$, leveraging the theoretical invertibility of the unconditional PF-ODE trajectory defined in Eq. 11. However, since both $\boldsymbol{\Phi}$ and $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}$ are both implemented in a first-order ODE solver in this paper, perfect invertibility cannot be guaranteed. Nevertheless, we find that this approximation strategy not only eliminates the need to compute the U-Net Jacobian—reducing computational cost—but also improves optimization stability, as demonstrated in Sec. 6.
4.5 Theoretical Analysis of Distillation Error
We theoretically prove that our SCTD loss leads to a tighter upper bound on the distillation error, explaining why SCTD achieves superior generative results. Specifically, the SDS loss can be reformulated as:
$\mathcal{L}_{\text{SDS}}$ $=\mathbb{E}_{t}[\omega(t)||\boldsymbol{\epsilon_{\phi}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})-\boldsymbol{\epsilon}^{*}||_{2}^{2}]$ (12)
$=\mathbb{E}_{t}[b(t)||\mathbf{z}_{s_{m}}-\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})||_{2}^{2}].$
Based on Eq. 12, we have the following theorem hold:
Theorem 1.
Given a sub-trajectory $[s_{m},s_{m+1}]$, let $\Delta t=max_{t,s\in(s_{m},s_{m+1}]}\{|t-s|\}$. We assume $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}$ satisfies the Lipschitz condition. If there exists a $\mathbf{z}_{0}$ satisfying $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})=\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})$ for $\forall t,s\in[s_{m},s_{m+1}]$ and a true image $\mathbf{z}^{data}\sim\mathbb{P}_{data}(\mathbf{z})$ corresponding to $\mathbf{z}_{0}$, we have
$\sup_{t,s\in[s_{m},s_{m+1})}||\mathbf{z}_{0}-\mathbf{z}^{data}||\\ =\mathcal{O}(\Delta t)(s_{m+1}-s_{m}),$ (13)
Proof.
The full proof, inspired by consistency model-related works [27, 44, 48, 22], is provided in App. E. ∎
Theorem 1 shows that our SCTD loss achieves an arbitrarily small distillation error when the ODE solver’s step size $\Delta_{t}$ and the segment length $s_{m+1}-s_{m}$ are sufficiently small. By choosing an appropriate segment number $N_{s}$, we obtain a tighter error upper bound, improving upon the bound $\mathcal{O}(\Delta t)(T)$ in CDS [48] and $\mathcal{O}(\Delta t)(T-e)$ in GCS [22]. This refinement enables more accurate gradient for optimizing the 3D representation $\boldsymbol{\theta}$, leading to higher-quality generative results.
5 Experiments
Figure 4: Qualitative comparisons with two SD-based metheds: DreamFusion [31] and LucidDreamer [23], and two CD-based methods: Consistent3D [48] and ConnectCD [22]. CFG scales are set to 100, 7.5, 20$\sim$40, 7.5, 7.5, respectively. Our approach yields high-quality results within the shortest time. Please zoom in for details.
5.1 Implementation Details
Our text-to-3D algorithm is implemented in PyTorch and trained on an A100 GPU for 5,000 iterations using the Adam optimizer [19]. We utilize Stable Diffusion 2.1 [34] as our foundational generative model and employ Point-E [30] to initialize 3D Gaussians. Hyperparameters related to 3D Gaussian Splatting (3DGS), such as learning rates and camera poses, are set following the guidelines in previous works [23, 22]. $t\sim\mathcal{U}(20,500+t_{warm})$ where $t_{warm}$ decreases linearly from 480 to 0 during the first 1,500 epochs.
5.2 Text-to-3D Generation
Qualitative Comparison. We present a collection of 3D assets generated by SegmentDreamer from different camera views in Fig. 1. Our algorithm effectively produces high-quality, text-aligned 3D assets. To highlight the superior generative capability of SegmentDreamer, we conduct qualitative experiments comparing it with state-of-the-art methods based on Consistent Distillation (CD)—namely, Consistent3D [48] and Connect3D [22]—as well as Score Distillation (SD)-based methods, including DreamFusion [31] and LucidDreamer [23], as illustrated in Fig. 4. For a fair comparison, all methods use Stable Diffusion 2.1 as the foundational generative model. As shown in Fig. 4, Both DreamFusion and Consistent3D suffer from severe Janus artifacts and fail to generate accurate conditional outputs. LucidDreamer encounters issues with overexposure and ineffective conditional generation. Connect3D performs well in both conditional generation and color rendering; however, its results still contain undesired artifacts, particularly in the tiger and beagle models (see Fig. 4). Moreover, as a 3DGS-based method, it requires over an hour to generate a single 3D asset, limiting its practical usability. In contrast, SegmentDreamer can generate high-quality results within about 32 minutes using CFG [9] and 38 minutes using Perp-Neg [1], proving its efficiency and superiority. More comparison results are shown in our supplementary material.
We also compare our SCTD loss with existing CD-based losses, CDS [48] and GCS [22]. For a fair comparison, we use Stable Diffusion 2.1 as the generative model and keep the random seed, guidance scale, and 3DGS-related parameters consistent across all methods. As shown in Fig. 5, CDS struggles to provide effective conditional guidance at a normal CFG scale, as also noted in [22]. Even at a higher CFG scale, its results remain unsatisfactory due to the lack of cross-consistency in its distillation loss. GCS, meanwhile, suffers from overexposure and artifacts due to excessive conditional guidance. While it attempts to address this with Brightness-Equalized Generation (BEG), artifacts persist (see Fig. 5, zoom in for details). In contrast, our SCTD mitigates ineffective conditional generation, overexposure, and artifacts, striking a better balance between self-consistency and cross-consistency.
Figure 5: Visual comparisons with CDS [48], GCS [22], and GCS+BEG [22]. Artifacts are marked with red circles.
Figure 6: Ablation study of the trajectory segmentation strategy with segment numer $N_{s}$ and time theshold $t_{\tau}$. Prompt: a DSLR photo of the Imperial State Crown of England.
Figure 7: Visual comparison results before and after performing the approximation strategy proposed in Sec. 4.3. “w/ Approx.” and “w/o Approx.” indicate whether the approximation strategy is used or not. The GFG scale is set to 7.5 for all generations.
Quantitative Comparison. We assess text alignment and human preference for each text-to-3D method in Tab. 1, where we generate 3D models from 40 prompts selected from the DreamFusion Gallery  [^1] https://dreamfusion3d.github.io/gallery.html and compute the average CLIP Score and ImageReward (IR) score across 120 camera views per model. Following Connect3D, we compute the FID score between 40 objects (each renders 120 views) and 50,000 images generated by Stable Diffusion 2.1. As shown, our method achieves the best text alignment, human preference, and fidelity while requiring the least time, showing its superior performance. The 40 prompts are given in the supplementary material.
Table 1: Quantitative comparison using CLIP score, IR score, training time (min), FID score, and three user study questions given by 40 volunteers. Bold and underline texts denote the best and second-best performance, respectively.
[…]
DreamFusion[31] 28.47 -0.004 60 140.84 4.73 4.87 4.90
LucidDreamer[23] 29.99 0.006 45 121.80 2.93 2.98 2.93
Consistent3D[48] 30.60 0.004 140 113.61 4.14 3.88 3.93
ConnectCD [22] 30.73 0.018 80 112.61 1.63 1.80 1.93
Ours 30.88 0.020 38 110.45 1.57 1.47 1.33
User Study. We conduct a user study with 40 volunteers to assess the generation quality for each algorithm. Each volunteer reviews eight groups of five 360-degree rendered videos and ranks them from the perspective of text alignment (Q1), object realism (Q2), and object details (Q3, e.g., artifact, brightness, Janus problem). To minimize bias, video order is randomized. As shown in Tab. 1, users consistently favor our method across all criteria.
6 Ablation Study
Effect of Trajectory Segmentation Strategy We visualize the effect of different trajectory segmentation strategies applied to our algorithm in Fig. 6. When the segment number $N_{s}$ and the time threshold $t_{\tau}$ are fixed, we observe that the segmentation strategy has minimal impact on the generated results. This is primarily because the segmentation strategy essentially determines $\sup|t-e|$. However, the timesteps $t$ and $s$ are randomly selected, $|t-s_{m}|$, $|t-s|$, and $|s-s_{m}|$ vary randomly during each training iteration, with their values potentially being large or small.
Effect of Segment Number $N_{s}$. We visualize the effect of different segment numbers $N_{s}$ in Fig. 6. As observed, a small $N_{s}$, regardless of the segmentation strategy, consistently results in overly smoothed outputs. As $N_{s}$ increases, the generation quality improves. However, when $N_{s}$ becomes too large, e.g. 10, while more details emerge, the overall model representation becomes vague, which is also mentioned in LucidDreamer [23]. We find that $N_{s}=5$ performs well in most cases.
Effect of time threshold $t_{\tau}$. As shown in Fig. 6, a large $t_{\tau}$ also leads to overly smoothed results for both segmentation strategies. This may be because a limited number of deterministic sampling steps in the diffusion process fails to effectively preserve the information of the initiation point $\mathbf{z}_{0}$, highlighting the necessity of incorporating multi-step deterministic sampling.
Validation of Approximation Strategy. We analyze the impact of replacing $\boldsymbol{G}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$ with $\mathbf{z}_{s_{m}}$ in Fig. 7. As shown, $\boldsymbol{G}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$ introduces slight overexposure and geometry distortion in generation results. In contrast, $\mathbf{z}_{s_{m}}$ not only produces better generation results but also reduces the training time. This suggests a gap between theory and practice, resembling the omission of U-Net Jacobian in DreamFusion [31].
7 Conclusion
We analyze improper conditional guidance in previous Consistency Distillation methods and identify its root cause as an imbalance between self-consistency and cross-consistency in denoising trajectories. Additionally, we find that large distillation error upper bounds limit generative capability. To address this, we propose the Segmented Consistency Trajectory Distillation loss, which improves conditional guidance and tightens the distillation bound. Experiments show that our SegmentDreamer achieves high-quality generation efficiently, advancing distillation-based text-to-3D methods toward practical use.
Limitation. SegmentDreamer mainly focuses on single-instance generation and performs suboptimally in multi-instance generation.
Potential Negative Impact. SegmentDreamer focuses on content systhesis, which may be used to create misleading content.
Supplementary Material
To provide proofs of SegmenDreamer and visual comparisons of state-of-the-arts, this supplementary material includes the following contents:
  • Section A: Guided Consistency Sampling Loss
  • Section B: Flaws in the Consistency Function of Guided Consistency Sampling
  • Section C: How to Connect SCTD with SDS
  • Section D: Computation Reduction Trick
  • Section E: Proof of Upper Bound of Distillation Error
  • Section F: Visual comparisons of State-of-the-art
  • Section G: Limitation and Potential Negative Impacts
A Guided Consistency Sampling Loss
The Guided Consistency Sampling (GCS) loss [22] is composed of a compact consistency loss $\mathcal{L}_{\text{CC}}$, a conditional guidance loss $\mathcal{L}_{\text{CG}}$, and a pixel-wise constraint loss $\mathcal{L}_{\text{CP}}$, which are defined by
$\mathcal{L}_{\text{CC}}(\boldsymbol{\theta})=\mathbb{E}[\left\|\boldsymbol{G_{\theta}}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,e,\emptyset)-\boldsymbol{G_{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,e,\emptyset)\right\|_{2}^{2}],$ (14)
$\mathcal{L}_{\text{CG}}(\boldsymbol{\theta})=\mathbb{E}[\left\|\boldsymbol{F_{\theta}}(\mathbf{z}_{e},e,\emptyset)-\boldsymbol{F_{\theta}}(\boldsymbol{G}_{\theta}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,e,\mathbf{y}),e,\emptyset)\right\|_{2}^{2}],$
$\mathcal{L}_{\text{CP}}(\boldsymbol{\theta})\!=\!\mathbb{E}[\left\|\!\boldsymbol{D}(\!F_{\theta}(\mathbf{z}_{e},e,\emptyset))\!\!-\!\!\boldsymbol{D}(\boldsymbol{F_{\theta}}(\boldsymbol{G}_{\theta}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,e,\mathbf{y}),e,\mathbf{y})\!)\!\right\|_{2}^{2}],$
where $e<s<t$, $\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t}$ is estimated by the following trajectory: $\mathbf{z}_{e}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}^{*}\to\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}=\boldsymbol{\Phi}(\mathbf{z}_{e},e,s,\emptyset)\to\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t}=\boldsymbol{\Phi}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{s},s,t,\emptyset)$, $\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}=\boldsymbol{\Phi}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,s,\mathbf{y})$, and $\boldsymbol{D}$ denotes the VAE decoder.
B Flaws in the Consistency Function of Guided Consistency Sampling
As we know, given a well-trained diffusion model $\boldsymbol{\phi}$, there exists an exact solution from timestep $t$ to $e$ [26]:
$\boldsymbol{G}(\mathbf{z}_{t},t,e,\mathbf{y})\!=\!\frac{\alpha_{e}}{\alpha_{t}}\mathbf{z}_{t}+\alpha_{s}\!\int_{\lambda_{t}}^{\lambda_{e}}e^{-\lambda}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t_{\lambda}(\lambda)},t_{\lambda}(\lambda),\mathbf{y})\mathrm{d}\lambda,$ (15)
where $\lambda_{t}=\mathrm{ln}\frac{\alpha_{t}}{\sigma_{t}}$ and $t_{\lambda}$ denotes the inverse function of $\lambda_{t}$. Inspired by [44], we find Eq. 6 suggests that GCS aims to optimize a 3D representation $\boldsymbol{\theta}$ such that $\boldsymbol{G_{\theta}}(\mathbf{z}_{t},t,e,\emptyset)=\boldsymbol{G_{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,e,\emptyset)$ for $\forall t,s,e\in[0,T]$ where $t>s>e$. This implies $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})$, as defined in Eq. 5, is not an approximation but an exact solution learning related to the 3D representation $\boldsymbol{\theta}$, i.e., $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\int_{\lambda_{t}}^{\lambda_{e}}e^{-\lambda}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t_{\lambda}(\lambda)}^{\mathbf{c}},t_{\lambda}(\lambda),\mathbf{y})\mathrm{d}\lambda}{\int_{\lambda_{t}}^{\lambda_{e}}e^{-\lambda}\mathrm{d}\lambda}$. However, dropping the target timestep $e$ in $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})$, as GCS does, is problematic. Suppose we predict both $\mathbf{z}_{e}$ and $\mathbf{z}_{e^{\prime}}$ from $\mathbf{z}_{t}$ where $t>e^{\prime}>e$, we must have
$\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\int_{\lambda_{t}}^{\lambda_{e^{\prime}}}e^{-\lambda}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t_{\lambda}(\lambda)},t_{\lambda}(\lambda),\mathbf{y})\mathrm{d}\lambda}{\int_{\lambda_{t}}^{\lambda_{e^{\prime}}}e^{-\lambda}\mathrm{d}\lambda}$ (16)
$\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\int_{\lambda_{t}}^{\lambda_{e}}e^{-\lambda}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t_{\lambda}(\lambda)},t_{\lambda}(\lambda),\mathbf{y})\mathrm{d}\lambda}{\int_{\lambda_{t}}^{\lambda_{e}}e^{-\lambda}\mathrm{d}\lambda}.$
Clearly, without the target timestep in $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})$, optimizing a 3D representation $\boldsymbol{\theta}$ to satisfy both conditions in Eq. 16 for all intervals $[e,t]$ is invalid. For GCS, as the number of training steps increases, the above unreasonable phenomena occur frequently, potentially resulting in poor distillation results.
C How to Connect SCTD with SDS?
As described in Sec. 4.3, given any subtrajectory $[s_{m},s_{m+1}]$, $\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,s_{m},\mathbf{y}):=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\alpha_{s_{m}}}{\alpha_{t}}\mathbf{z}_{t}-\alpha_{s_{m}}\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda$. Then, we have
$\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\mathbf{z}_{t},t,\mathbf{y})-\frac{\alpha_{s_{m}}}{\alpha_{t}}\mathbf{z}_{t}}{\alpha_{s_{m}}\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda},$ (17)
where $\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$. In this case, $\mathbf{y}$ can represent any prompt embedding, including $\emptyset$. According to Eq. 7, $\mathcal{L}_{\text{SDS}}$ can be further transformed into:
$\hat{\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}=\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})+\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})$ (18)
$+\omega(\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\emptyset)),$
where $\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\mathbf{z}_{t},t,s,\mathbf{y})$. By substituting Eq. 17 into Eq. 18, we obtain
$\mathcal{L}_{\text{SDS}}(\boldsymbol{\theta})\!=\!\mathbb{E}_{t}[b(t)\!||\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\mathbf{y})$ (19)
$+\omega(\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\emptyset)\!-\!\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\mathbf{y}))\!-\!\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})$
$+\frac{\alpha_{s_{m}}}{\alpha_{t}}\mathbf{z}_{t}-\alpha_{s_{m}}\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda\boldsymbol{\epsilon}||_{2}^{2}],$
where $b(t)=\frac{\omega(t)}{(\alpha_{s_{m}}\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda)^{2}}$. Furthermore,
$\frac{\alpha_{s_{m}}}{\alpha_{t}}\mathbf{z}_{t}-\alpha_{s_{m}}\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda\boldsymbol{\epsilon}$ (20)
$=\frac{\alpha_{s_{m}}}{\alpha_{t}}(\alpha_{t}\mathbf{z}_{0}^{\mathbf{c}}+\sigma_{t}\boldsymbol{\epsilon})+\alpha_{s_{m}}(e^{\lambda_{s_{m}}-\lambda_{t}}-1)\boldsymbol{\epsilon}$
$=\alpha_{s_{m}}\mathbf{z}_{0}+\sigma_{s_{m}}\boldsymbol{\epsilon}$
$=\mathbf{z}_{s_{m}}$
Substituting Eq. 20 into Eq. 19, we can obtain Eq. 8. Proof is completed.
D Computation Reduction Trick
Eq. 7 can be transformed into
$\hat{\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})\!-\!\boldsymbol{\epsilon}\!=\!\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})\!-\!\boldsymbol{\epsilon}\!+\!\omega(\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})\!-\!\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\emptyset))$ (21)
$=\!\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\emptyset)\!-\!\boldsymbol{\epsilon}\!+\!(\omega+1)(\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})\!-\!\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\emptyset))$
Based on Eq. 21, we can readily derive Eq. 9 by following the procedure described in App. C.
E Proof of Upper Bound of Distillation Error
Before proving Theorem 1, we first give the following lemma:
Lemma 1.
Given a sub-trajectory $[s_{m},s_{m+1}]$, let $\Delta t=max_{t,s\in[s_{m},s_{m+1})}\{|t-s|\}$. We assume $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}$ satisfies the Lipschitz condition and the ODE solver has local error uniformally bounded by $O(t-s)^{p+1}$ with $p\geq 1$. If $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\emptyset)=\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\emptyset)$ for $\forall t,s\in[s_{m},s_{m+1}]$, we have
$\sup_{t,s\in[s_{m},s_{m+1})}\!\!\!\!\!||\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,\mathbf{y})-\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,s_{m},\mathbf{y})||$ (22)
$=O((\Delta t)^{p})(s_{m+1}-s_{m}).$
Proof.
The proof is based on [40, 44, 22, 48]. Let
$\boldsymbol{e}_{n-1}:=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,s_{m},\mathbf{y}),$ (23)
where $\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\mathbf{z}_{s_{m}}^{\mathbf{c}},s_{m},s,\mathbf{y})$. According to the condition, we have
$\boldsymbol{e}_{n}=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})-\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,s_{m},\mathbf{y})$ (24)
$=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})$
$+\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,s_{m},\mathbf{y})$
$=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})+\boldsymbol{e}_{n-1},$
Provided that $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}$ satisfies $L$-Lipschitz condition, we have
$||e_{n}||=||e_{n-1}+\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})||$ (25)
$\leq||e_{n-1}||+||\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})||$
$\leq\boldsymbol{e}_{n-1}+L||\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}-\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}||$
$\leq\boldsymbol{e}_{n-1}+L\cdot O((t-s)^{p+1})$
$\leq\boldsymbol{e}_{n-1}+L(t-s)\cdot O((\Delta t)^{p}).$
Besides, according to the boundray condition,
$\boldsymbol{e}_{s_{m}}=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\hat{\mathbf{z}}_{s_{m}}^{\boldsymbol{\Phi}},s_{m},\mathbf{y})-\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{s_{m}}^{\boldsymbol{\Phi}},s_{m},s_{m},\mathbf{y})$ (26)
$=\hat{\mathbf{z}}_{s_{m}}^{\boldsymbol{\Phi}}-\tilde{\mathbf{z}}_{s_{m}}^{\boldsymbol{\Phi}}=\mathbf{z}_{s_{m}}^{\mathbf{c}}-\mathbf{z}_{s_{m}}^{\mathbf{c}}=0,$
Therefore,
$||\boldsymbol{e}_{n}||\leq||\boldsymbol{e}_{s_{m}}||+L\sum_{t_{i},t_{i-1}\in[s_{m},s_{m+1}]}(t_{i}-t_{i-1})O((\Delta t)^{p})$ (27)
$=O((\Delta t)^{p})\cdot(s_{m+1}-s_{m}).$
The proof is completed. ∎
According to Eq. 12, one can ideally optimize a 3D model $\boldsymbol{\theta}$ such that $\mathbf{z}_{s_{m}}=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$. In this case, we have $\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,s_{m},\mathbf{y})=\mathbf{z}_{s_{m}}^{data}$, where $\mathbf{z}_{s_{m}}^{data}=\alpha_{s_{m}}\mathbf{z}^{data}+\sigma_{s_{m}}\boldsymbol{\epsilon}^{*}$. Based on this, we have
$||\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,\mathbf{y})-\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,s_{m},\mathbf{y})||\!=\!||\mathbf{z}_{s_{m}}\!-\!\mathbf{z}^{data}_{s_{m}}||\!=\!||\mathbf{z}_{0}\!-\!\mathbf{z}^{data}||.$ (28)
Since we use a first-order ODE solver to implement $\boldsymbol{\Phi}$, we have
$\sup_{t,s\in[s_{m},s_{m+1})}||\mathbf{z}_{0}-\mathbf{z}^{data}||\\ =\mathcal{O}(\Delta t)(s_{m+1}-s_{m}).$ (29)
The proof is completed.
F Visual comparisons of State-of-the-arts
We also provide visual comparisons of State-of-the-arts, as shown in Fig. 8. We provide additional qualitative comparisons with DreamFusion [31], LucidDreamer [23], Consistent3D [48], and ConnectCD [22]. CFG scales are set to 100, 7.5, 20$\sim$40, 7.5, 7.5, respectively. Our approach yields results with high quality within the shortest time. In Table 2, we provide the detailed prompts of examples.
G Limitation and Potential Negative Impact
Limitation. SegmentDreamer is primarily designed for single-instance generation and performs suboptimally in multi-instance generation.
Potential Negative Impact. SegmentDreamer focuses on content synthesis, which could potentially be misused to create misleading or deceptive content.
Figure 8: Additional qualitative comparisons with DreamFusion [31], LucidDreamer [23], Consistent3D [48], and ConnectCD [22]. CFG scales are set to 100, 7.5, 20$\sim$40, 7.5, 7.5, respectively. Our approach yields results with high quality within the shortest time. Please zoom in for details.
Table 2: Prompt of examples.
Prompt
1 A cat with a mullet
2 A pig wearing a backpack
3 A DSLR photo of an origami crane
4 A photo of a mouse playing the tuba
5 An orange road bike
6 A ripe strawberry
7 A DSLR photo of the Imperial State Crown of England
8 A photo of a wizard raccoon casting a spell
9 A DSLR photo of a corgi wearing a top hat
10 A rabbit, animated movie character, high-detail 3D model
11 A panda rowing a boat
12 A highly detailed sand castle
13 A DSLR photo of a chimpanzee dressed like Henry VIII king of England
14 A photo of a skiing penguin wearing a puffy jacket, highly realistic DSLR photo
15 A blue poison-dart frog sitting on a water lily
16 A DSLR photo of a bear dressed in medieval armor
17 A DSLR photo of a squirrel dressed like a clown
18 A plush toy of a corgi nurse
19 A humanoid robot playing the violin
20 A DSLR photo of a bear dressed as a lumberjack
21 A blue motorcycle
22 Michelangelo style statue of an astronaut
23 A DSLR photo of a chow chow puppy
24 A DSLR photo of cats wearing eyeglasses
25 A red panda
26 A DSLR photo of an elephant skull
27 An amigurumi bulldozer
28 A typewriter
29 A red-eyed tree frog, low poly
30 A DSLR photo of a chimpanzee wearing headphones
31 A robot made out of vegetables
32 A DSLR photo of a red rotary telephone
33 A DSLR photo of a blue lobster
34 A DSLR photo of a squirrel flying a biplane
35 A DSLR photo of a baby dragon hatching out of a stone egg
36 A DSLR photo of a bear dancing ballet
37 A plate of delicious tacos
38 A DSLR photo of a car made out of cheese
39 A yellow school bus
40 A DSLR photo of a shiny beetle