### v2

# SegmentDreamer: Towards High-fidelity Text-to-3D Synthesis with Segmented Consistency Trajectory Distillation

Abstract
Recent advancements in text-to-3D generation improve the visual quality of Score Distillation Sampling (SDS) and its variants by directly connecting Consistency Distillation (CD) to score distillation. However, due to the imbalance between self-consistency and cross-consistency, these CD-based methods inherently suffer from improper conditional guidance, leading to sub-optimal generation results. To address this issue, we present SegmentDreamer, a novel framework designed to fully unleash the potential of consistency models for high-fidelity text-to-3D generation. Specifically, we reformulate SDS through the proposed Segmented Consistency Trajectory Distillation (SCTD), effectively mitigating the imbalance issues by explicitly defining the relationship between self- and cross-consistency. Moreover, SCTD partitions the Probability Flow Ordinary Differential Equation (PF-ODE) trajectory into multiple sub-trajectories and ensures consistency within each segment, which can theoretically provide a significantly tighter upper bound on distillation error. Additionally, we propose a distillation pipeline for a more swift and stable generation. Extensive experiments demonstrate that our SegmentDreamer outperforms state-of-the-art methods in visual quality, enabling high-fidelity 3D asset creation through 3D Gaussian Splatting (3DGS).
Figure 1: Examples of 3D assets generated by SegmentDreamer. Our framework addresses the improper conditional guidance issues of recent Consistency Distillation (CD)-based methods and theoretically provides a significantly tighter upper bound on distillation error, enabling high-fidelity text-to-3D generation within a reasonable timeframe ($\sim$32 minutes with classifier-free guidance [9] and $\sim$38 minutes combined with Perp-Neg [1]) on a single A100 GPU through 3D Gaussian Splatting (3DGS). Project Page: https://zjhJOJO.github.io/segmentdreamer
footnotetext: *Corresponding author.
1 Introduction
Text-to-3D generation aims to create 3D assets that accurately align with text descriptions. It plays a vital role in various domains such as virtual reality, 3D gaming, and film production. One compelling solution is to directly “distill” 3D assets from 2D text-to-image Diffusion Models (DM) [34] through Score Distillation Sampling (SDS) [31, 46], which leverages pre-trained text-to-image DMs as the 3D priors to guide the optimization of differentiable 3D representations, such as Neural Radiance Fields (NeRF) [29] and 3D Gaussian Splatting (3DGS) [17]. Subsequent research focuses on improving the performance of SDS in diversity [47, 11, 43], visual quality [3, 24, 47, 51, 58, 48, 22, 25, 53, 57, 16], and multi-view consistency [5, 10, 37, 35, 12].
Figure 2: Visual examples of existing CD-based methods: CDS [48], GCS [22], and our SegmentDreamer with a Classifier-Free Guidance (CFG) scale of 7.5. As shown, improper conditional guidance leads to suboptimal results. (a) Ineffective guidance produces semantically inconsistent details, while (b) excessive guidance results in overexposure and artifacts. Although (c) Brightness-Equalized Generation (BEG) [22] can mitigate overexposure, artifacts persist. In contrast, our SegmentDreamer enhances conditional generation, producing high-fidelity 3D assets.
Recent methods [48, 22] have improved SDS and its variants by incorporating Consistency Distillation (CD) [40, 18] into their distillation loss. However, due to an imbalance between self-consistency and cross-consistency in their losses, these methods inherently suffer from improper conditional guidance when combined with Classifier-Free Guidance (CFG) for generation, ultimately leading to suboptimal outcomes (see Figure 2; a detailed explanation is provided in Section 4.1). Additionally, the distillation error upper bound of these methods remains large, limiting their generation quality.
To address the aforementioned limitations, we present a novel framework – SegmentDreamer, which aims to fully unleash the potential of consistency models for high-fidelity text-to-3D generation. Specifically, by analyzing existing CD-based methods, we draw inspiration from segmented consistency models [33, 44] to reformulate SDS through the proposed Segmented Consistency Trajectory Distillation (SCTD), which effectively mitigates the imbalance between self-consistency and cross-consistency by explicitly defining their relationship. Moreover, unlike CD-based methods that directly enforce consistency across the entire Probability Flow Ordinary Differential Equation (PF-ODE) trajectory, SCTD partitions the PF-ODE trajectory into multiple sub-trajectories, ensuring consistency within each segment. This approach theoretically provides a significantly tighter upper bound on distillation error. In addition, we propose a distillation pipeline for a more swift and stable generation. We also introduce 3DGS as our 3D representation. Extensive experimental results demonstrate that our SegmentDreamer outperforms state-of-the-art methods in visual quality. Moreover, the proposed SCTD can be seamlessly applied to various 3D generative tasks (e.g., 3D avatar generation and 3D portrait generation), facilitating the creation of high-fidelity 3D assets.
The main contributions are summarized as follows:
  • We present SegmentDreamer, a novel framework for high-fidelity text-to-3D generation, which addresses the imbalance issue of existing CD-based methods.
  • We reformulate SDS through the proposed SCTD, effectively mitigating the imbalance between self- and cross-consistency. SCTD also theoretically provides a significantly tighter upper bound on distillation error. We further propose a distillation pipeline for more swift and stable optimization.
  • Experimental results demonstrate that our SegmentDreamer outperforms state-of-the-art methods in visual quality, enabling high-fidelity 3D asset generation through 3DGS.
2 Related Work
2.1 Text-to-3D Generation
Creating 3D objects from text descriptions is challenging in the fields of computer vision and graphics. To achieve this goal, a series of end-to-end text-to-3D pipelines have been proposed [8, 7, 30, 13, 21, 49], but these methods require large-scale labeled 3D datasets for training. To reduce the reliance on 3D training data, a compelling solution is to directly “distill” 3D assets from 2D text-to-image Diffusion Models (DM) [34]. A groundbreaking contribution in this domain is DreamFusion [31], which introduced the Score Distillation Sampling (SDS) loss to optimize parametric 3D representations. Follow-up works focus on refining and extending SDS in various aspects. For example, several studies [47, 11, 43] improve SDS in diversity through prompting modifications, while others [3, 24, 47, 51, 58, 48, 22, 25, 53, 57, 16] enhance the fidelity of the generated 3D objects. Other advancements aim to achieve multi-view consistency [5, 10, 37, 35, 12] and faster generation [42, 52]. Recently, Consistency Distillation (CD) [40, 18] has received widespread attention as it enables high-quality generation through minimal sampling steps. Inspired by CD, Wu et al. [48] and Li et al. [22] incorporate CD into their distillation loss, which addresses some limitations brought by SDS. However, these CD-based methods inherently suffer from improper conditional guidance, leading to sub-optimal results. In this work, We aim to unleash CD’s potential in generating high-fidelity 3D assets and push text-to-3D synthesis boundaries.
2.2 Differentiable 3D Representations
Neural Radiance Field (NeRF) [29] aims to construct implicit 3D representations using a multi-layer perceptron (MLP), leveraging standard volumetric rendering [14] and alpha compositing techniques [32]. Subsequent research has explored adapting NeRF to various applications, including sparse-view reconstruction [45], anti-aliasing [2], and medical image super-resolution [4].
3D Gaussian Splatting (3DGS) [17] has emerged as a leading approach for 3D representation, offering rapid optimization and high-speed rendering. Unlike NeRF, 3DGS explicitly models scenes using a set of 3D Gaussians and renders views through splatting [20]. Its versatility has enabled applications across various domains, including avatar generation [36, 55], single-view reconstruction [41], anti-aliasing [54], digital watermarking [6], and SLAM [28, 50].
3 Preliminaries
Diffusion Model (DM) [39, 38, 15] consists of a diffusion process and a denoising process. The diffusion process can be modeled as the solution $\hat{\text{o}}$ stochastic differential equation (SDE) with the transition kernel $\mathbb{P}(\mathbf{z}_{t}|\mathbf{z}_{0}):=\mathcal{N}(\mathbf{z}_{t};\alpha_{t}\mathbf{z}_{0},\sigma_{t}^{2}\mathbf{I})$. By reversing the diffusion process, $\mathbf{z}_{0}$ can be estimated by solving a Probability Flow Ordinary Differentiable Equation (PF-ODE), which is defined by $\mathrm{d}\mathbf{z}_{t}=[f(t)\mathbf{z}_{t}-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{z}_{t}}\mathrm{log}\mathbb{P}(\mathbf{z}_{t})]\mathrm{d}t$. A neural network ${\boldsymbol{\epsilon_{\phi}}}(\mathbf{z},t)$ can be trained to approximate the scaled score function $-\sigma_{t}\nabla_{\mathbf{z}_{t}}\mathrm{log}\mathbb{P}(\mathbf{z}_{t})$, offering an empirical formulation of the PF-ODE: $\mathrm{d}\mathbf{z}_{t}=[f(t)\mathbf{z}_{t}+\frac{g(t)^{2}}{2\sigma_{t}}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t)]\mathrm{d}t.$
Consistency Model [40, 27] aims to achieve few-step inference by distilling a pre-trained DM $\boldsymbol{\phi}$ into a student model $\boldsymbol{\phi}_{stu}$. Given a PF-ODE trajectory $\{\mathbf{z}_{t}\}_{t\in[0,T]}$ and a condition $\mathbf{y}$, a consistency model $\boldsymbol{f}_{\boldsymbol{\phi}_{stu}}$ builds the mapping $\boldsymbol{f}_{\boldsymbol{\phi}_{stu}}(\mathbf{z}_{t},t,\mathbf{y})=\mathbf{z}_{0}$ between any points along this trajectory and the endpoint $\mathbf{z}_{0}$ by minimizing the following Consistency Distillation (CD) loss:
$\mathcal{L}_{\text{CD}}=\mathbb{E}[\lambda(t)||\boldsymbol{f}_{\boldsymbol{\phi}_{stu}}(\mathbf{z}_{t},t),\boldsymbol{f}_{\boldsymbol{\phi}_{stu}^{-}}(\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s},s)||_{2}^{2}],t>s,$ (1)
where $\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ for $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\boldsymbol{\phi}_{stu}^{-}$ is updated with the exponential moving average [40, 27]. $\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}$ can be obtained by:
$\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}$ $=\boldsymbol{\Phi}(\mathbf{z}_{t},t,s,\mathbf{y})$ (2)
$=\mathbf{z}_{t}+\int_{t}^{s}[f(u)\mathbf{z}_{u}+\frac{g(u)^{2}}{2\sigma_{u}}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{u},u,\mathbf{y})]\mathrm{d}u.$
The consistency function $\boldsymbol{f}_{\boldsymbol{\phi}_{stu}}$ can be parameterized in various ways [40, 27], and the solver $\boldsymbol{\Phi}(\cdot,\cdot,\cdot,\cdot)$ is used to calculate the point at $s$, which can be implemented using DDIM [38] or DPM-solver [26].
Score Distillation Sampling (SDS) Loss. Given a 3D representation $\boldsymbol{\theta}$, one can render a view image $\mathbf{z}_{0}:=\boldsymbol{g}(\mathbf{c},\boldsymbol{\theta})$ with camera pose $\mathbf{c}$, SDS distills a 3D object from a 2D DM $\boldsymbol{\phi}$ by minimizing the following SDS loss w.r.t. $\boldsymbol{\theta}$:
$\mathcal{L}_{\text{SDS}}:=\mathbb{E}_{t,\boldsymbol{\epsilon}}[\omega(t)||\underbrace{\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}}_{\text{noise residual}}||_{2}^{2}],$ (3)
where $\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$. Poole et al. [31] omit the Jacobian term $\frac{\partial\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})}{\partial\mathbf{z}_{t}}$ in gradient calculation for speedup.
4 Our Method
In this section, we first review the mechanism of existing Consistency Distillation (CD)-based methods [48, 22] and reveal that they inherently suffer from improper conditional guidance when combined with Classifier-free Guidance (CFG). Subsequently, we address these limitations by reformulating SDS through the lens of the proposed Segmented Consistency Trajectory Distillation (SCTD). Moreover, building on this insight, we propose an efficient SCTD sampling method for high-fidelity text-to-3D generation. We further tailor a distillation pipeline for the proposed SCTD sampling to achieve more swift and stable optimization. Finally, we provide a theoretical proof for our distillation error upper bound. The overall framework of our SegmentDreamer is depicted in Fig. 3, with the algorithm detailed in Algorithm 1.
4.1 Reviewing CD-based Losses
The consistency function w.r.t. the 3D representation $\boldsymbol{\theta}$ is parameterized as $\boldsymbol{F_{\theta}}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\mathbf{z}_{t}-\sigma_{t}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})}{\alpha_{t}}$ [48, 22]. Before proceeding, we clarify two key concepts: self-consistency and cross-consistency. The former ensures that all points on the same ODE trajectory consistently map to their original points, as in consistency models [27, 40, 18]. The latter enforces alignment between unconditional and conditional ODE trajectories [22].
Figure 3: An overview of SegmentDreamer: We begin by initializing a 3D representation $\boldsymbol{\theta}$ using a 3D generator, such as Point-E [30]. In each iteration, we randomly render a batch of camera views $\mathbf{z}_{0}$ from $\boldsymbol{\theta}$ and diffuse them into $\mathbf{z}_{s_{m}}$ with fixed noise $\boldsymbol{\epsilon}^{*}$. Next, we transform $\mathbf{z}_{s_{m}}$ into $\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t}$ using either one-step or two-step unconditional deterministic sampling. During the denoising process, we first estimate $\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}$ through one-step conditional deterministic sampling from $\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t}$. Subsequently, we compute two consistency functions and utilize them to derive the loss $\mathcal{L}_{\text{SCTD}}$, which is ultimately employed to optimize $\boldsymbol{\theta}$.
Consistency Distillation Sampling (CDS) loss [48] is defined as
$\mathcal{L}_{\text{CDS}}(\boldsymbol{\theta})=\mathbb{E}_{t,s}[c(t)||\boldsymbol{F_{\theta}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{F_{\theta}}(\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s},s,\mathbf{y})||_{2}^{2}],$ (4)
[…]
Guided Consistency Sampling (GCS) [22] leverages the DPM-Solver-1 [26] as its consistency function:
$\boldsymbol{G_{\theta}}(\mathbf{z}_{t},t,s,\mathbf{y})=\frac{\alpha_{s}}{\alpha_{t}}\mathbf{z}_{t}-\alpha_{s}\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})\int_{\lambda_{t}}^{\lambda_{s}}e^{-\lambda}\mathrm{d}\lambda,$ (5)
where $\lambda_{t}=\mathrm{ln}\frac{\alpha_{t}}{\sigma_{t}}$. Based on Eq. 5, GCS proposes a more compact consistency (CC) loss $\mathcal{L}_{\text{CC}}$ defined as
$\mathcal{L}_{\text{CC}}(\boldsymbol{\theta})\!=\!\mathbb{E}_{t,s,e}[\left\|\boldsymbol{G_{\theta}}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,e,\emptyset)\!-\!\boldsymbol{G_{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,e,\emptyset)\right\|_{2}^{2}],$ (6)
where $e$$<$$s$$<$$t$. GCS also employs a conditional guidance loss $\mathcal{L}_{\text{CG}}$ and a pixel-wise constraint $\mathcal{L}_{\text{CP}}$ to enforce cross-consistency (see Supp. A for details).
GCS ensures self-consistency within sub-trajectory, but it still suffers from improper conditional guidance due to two key issues: (1) $\boldsymbol{G_{\theta}}$ used to enforce cross-consistency is inherently flawed because $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})$ lacks a target time-step, as detailed in App. B. (2) GCS employs $\boldsymbol{F}_{\boldsymbol{\theta}}$ (implemented via one-step DDIM) to enforce cross-consistency throughout the entire ODE trajectory in both latent and pixel spaces. However, as reported in its paper, it suffers from unstable generation due to the excessive conditional guidance, leading to overexposure and artifacts (more experimental analysis is provided in App. Fig. 1). Furthermore, this approach results in a less stringent upper bound on the distillation error, expressed as $\mathcal{O}(\Delta_{t})(T-e)$, which constrains its generation quality.
Table 1: Differences between CDS, GCS, and SCTD.“traj.” denotes trajectory.
CDS [48] GCS [22] SCTD (Ours)
Self-consistency ✓(whole traj.) ✓ (sub traj.) ✓ (sub traj.)
[…]
Error bound $\mathcal{O}(\Delta_{t})T$ $\mathcal{O}(\Delta_{t})(T\!-\!e)$ $\mathcal{O}(\Delta_{t})(s_{m+1}\!-\!s_{m})$
Summary: The improper conditional guidance in CDS and GCS stems mainly from an imbalance between self-consistency and cross-consistency. Additionally, their reliance on $\boldsymbol{F_{\theta}}$ to enforce consistency across the entire trajectory exacerbates this issue. Tab. 1 clarifies the differences between CDS, GCS, and ours.
4.2 Reframing SDS through SCTD
We address the previous limitations by reframing SDS through the lens of SCTD. Specifically, we first define our consistency function as follows: as existing segmented consistency models [44, 33] do, we partition the entire time-step range $[0,T]$ into $N_{s}$ sub-intervals (i.e., $N_{s}+1$ edge points), represented by $\{[s_{m},s_{m+1})|m\in\{0,\cdots,N_{s}-1\},s_{0}=0,s_{N_{s}}=T\}$ (the segmentation strategy is discussed in Sec. 4.3). We define our consistency function $\boldsymbol{G_{\theta}}$ using Eq. 5, aiming to enforce consistency within each sub-trajectory, i.e., $\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,s_{m},\mathbf{y})=\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{s},s,s_{m},\mathbf{y})$ for $s_{m}\leq s<t\leq s_{m+1}$. For brevity, we abbreviate $\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,s_{m},\mathbf{y})$ as $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\mathbf{z}_{t},t,\mathbf{y})$. Notably, our consistency function $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}$ avoids the issues inherent in GCS, as analyzed in Sec. 4.1 and Supp. B, because at each training stage, the target time step $s_{m}$ is always uniquely determined.
When CFG [9] with scale $\omega$ is used, the noise residual term in Eq. 3 can be transformed into
$\hat{\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}$ $=\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}$ (7)
$+\omega(\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\emptyset)).$
Based on Eq. 7, $\mathcal{L}_{\text{SDS}}$ can be expressed in the form of SCTD:
$\mathcal{L}_{\text{SDS}}(\boldsymbol{\theta})=\mathbb{E}_{t,s}[b(t)||\underbrace{\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\mathbf{y})}_{\text{self-consistency constraint}}+$ (8)
$\omega\underbrace{(\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\emptyset)\!-\!\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\mathbf{y}))}_{\text{cross-consistency constraint}}\!+\!\underbrace{\mathbf{z}_{s_{m}}-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})}_{\text{generative prior}}||_{2}^{2}].$
where $b(t)=\frac{\omega(t)}{(\alpha_{s_{m}}\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda)^{2}}$ and $\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\mathbf{z}_{t},t,s,\mathbf{y})$. The derivation is given in App. D. The first term on the r.h.s. of Eq. 8 enforces self-consistency across the conditional denoising sub-trajectory, referred to as the ”self-consistency constraint.” The second term enforces cross-consistency between the unconditional and conditional denoising sub-trajectories, termed the ”cross-consistency constraint.” The third term adheres to the naming conventions in CSD [53]. As observed, Eq. 8 explicitly defines the relationship between the self-consistency constraint and the cross-consistency constraint. Furthermore, as demonstrated in Sec. 4.3, SCTD provides a tighter upper bound on the distillation error.
4.3 SCTD Sampling
Sampling Design. A natural question arises: according to Eq. 8, does minimizing $\mathcal{L}_{SDS}$ necessarily enforce both self- and cross-consistency? The answer is no. This is because one cannot necessarily achieve $x_{1}=x_{2}$ and $y_{1}=y_{2}$ when $||x_{1}-x_{2}+\omega(y_{1}-y_{2})||_{2}^{2}=0$. Additionally, we observe that the condition $\mathbf{y}$ appears in all three loss terms in Eq. 8, which increases the computational cost, particularly when using Perp-Neg [1] or ControlNet [56].
To improve the computational efficiency, we perform the following equivalent transformation on Eq. 8:
$\mathcal{L}_{\text{SDS}}(\boldsymbol{\theta})=\mathbb{E}_{t,s}[b(t)||\underbrace{\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\emptyset)-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\emptyset)}_{\text{self-consistency constraint}}+$ (9)
$(\omega\!+\!1)\!\underbrace{(\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\emptyset)\!-\!\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\mathbf{y}))}_{\text{cross-consistency constraint}}\!+\!\underbrace{\mathbf{z}_{s_{m}}\!-\!\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\emptyset)}_{\text{generative prior}}\!||_{2}^{2}].$
The derivation is given in Supp. E.
To enforce self- and cross-consistency, we follow the suggestion from CSD [53] to omit the generative prior from Eq. 9, and then impose a stricter constraint on Eq. 8:
$\mathcal{L}_{\text{SCTD}}(\boldsymbol{\theta})=\mathbb{E}_{t,s}[b(t)[||\underbrace{sg(\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\emptyset))-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\emptyset)}_{\text{self-consistency constraint}}||_{2}^{2}$ (10)
$+||(\omega+1)\underbrace{(\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\emptyset)-sg(\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})))}_{\text{cross-consistency constraint}}||_{2}^{2}]].$
where $\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}}$ is obtained through multi-step deterministic sampling from $\mathbf{z}_{0}$, as discussed in Sec. 4.4. $\mathbf{z}_{s_{m}}=\alpha_{s_{m}}\mathbf{z}_{0}+\sigma_{s_{m}}\boldsymbol{\epsilon}^{*}$, with $\boldsymbol{\epsilon}^{*}$ fixed throughout the entire training. $\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,s,\mathbf{y})$, where $\boldsymbol{\Phi}$ is implemented by a first-order ODE solver in this work. $sg(\cdot)$ represents the stopping of gradient backpropagation.
Algorithm 1 Text-to-3D Generation via SCTD Loss
1:Initialization: threshold $t_{\tau}$, segmentation number $N_{s}$, and target prompt $\mathbf{y}$
2:Initialize $\boldsymbol{\theta}$ with a pre-trained 3D generator
3:while $\boldsymbol{\theta}$ is not converged do
4:  Sample $\mathbf{z}_{0}=\boldsymbol{g}(\mathbf{c},\boldsymbol{\theta})$ and $t\sim\mathcal{U}(0,T)$
5:  Obtain $s_{m}\!=\!max\{s_{i}|s_{i}\leq t,i=0,\cdots,N_{s}-1\}$
6:  Obtain $\mathbf{z}_{s_{m}}=\alpha_{s_{m}}\mathbf{z}_{0}+\sigma_{s_{m}}\boldsymbol{\epsilon}^{*}$
7:  Sample $s\sim\mathcal{U}(s_{m},t)$
8:  Predict $\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}}$ with Eq. 11 and $\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}\!=\!\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},t,s,\mathbf{y})$
9:  Obtain $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\emptyset)$ and $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$ with Eq. 5
10:  Obtain $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{SCTD}}$ with Eq. 10 and update $\boldsymbol{\theta}$.
Trajectory Segmentation Strategy. We propose two heuristic trajectory segmentation strategies: equal segmentation and monotonically increasing segmentation. In the first strategy, the entire time-step range $[0,T]$ is uniformly divided into $N_{s}$ segments, i.e., $s_{m+1}-s_{m}=\frac{T}{N_{s}}$. In the second strategy, we adopt a monotonically increasing interval segmentation method, where the $N_{s}$ intervals satisfy that $s_{m+1}-s_{m}=t_{\tau}+\frac{2m(T-N_{s}t_{\tau})}{N_{s}(N_{s}-1)}$ for $\in\{0,\cdots,N_{s}-1\}$. A comparison of the two strategies is presented in Tab. 2.
4.4 Swift and Stable Optimization Pipeline
In previous work [23, 48, 22], $\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}}$ is derived from deterministic sampling with a fixed number of steps, lacking flexibility. In this paper, we adopt the following dynamic adjustment strategy: .
$\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}}=\begin{cases}\boldsymbol{\Phi}(\boldsymbol{\Phi}(\mathbf{z}_{s_{m}},s_{m},s,\emptyset),s_{m},t,\emptyset),&\text{if }t>t_{\tau}\\ \boldsymbol{\Phi}(\mathbf{z}_{s_{m}},s_{m},t,\emptyset),&\text{if }t\leq t_{\tau}\end{cases},$ (11)
This strategy enables us to make a trade-off between generation quality and optimization speed, as evidenced in Tab. 2.
Consistency function Approximation. For faster optimization, we propose approximating $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\emptyset)$ as $\mathbf{z}_{s_{m}}$, leveraging the theoretical invertibility of the unconditional PF-ODE trajectory defined in Eq. 11. However, since both $\boldsymbol{\Phi}$ and $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}$ are implemented using a first-order ODE solver in this paper, perfect invertibility cannot be guaranteed. Nevertheless, we find that this approximation strategy not only eliminates the need to compute the U-Net Jacobian—reducing computational cost—but also improves optimization stability, as demonstrated in Tab. 2.
4.5 Theoretical Analysis of Distillation Error
We theoretically prove that our SCTD loss leads to a tighter upper bound on the distillation error, which explains why SCTD achieves superior generative results. Specifically, the SDS loss can be reformulated as:
$\mathcal{L}_{\text{SDS}}$ $=\mathbb{E}_{t}[\omega(t)||\boldsymbol{\epsilon_{\phi}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})-\boldsymbol{\epsilon}^{*}||_{2}^{2}]$ (12)
$=\mathbb{E}_{t}[b(t)||\mathbf{z}_{s_{m}}-\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})||_{2}^{2}].$
Based on Eq. 12, we have the following theorem:
Theorem 1.
Given a sub-trajectory $[s_{m},s_{m+1}]$, let $\Delta t=max_{t,s\in(s_{m},s_{m+1}]}\{|t-s|\}$. We assume $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}$ satisfies the Lipschitz condition. If there exists a $\mathbf{z}_{0}$ satisfying $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})=\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})$ for $\forall t,s\in[s_{m},s_{m+1}]$ and a true image $\mathbf{z}^{data}\sim\mathbb{P}_{data}(\mathbf{z})$ (corresponding to $\mathbf{z}_{0}$), we have
$\sup_{t,s\in[s_{m},s_{m+1})}||\mathbf{z}_{0}-\mathbf{z}^{data}||\\ =\mathcal{O}(\Delta t)(s_{m+1}-s_{m}).$ (13)
Proof.
The full proof, inspired by consistency model-related works [27, 44, 48, 22], is provided in App. F. ∎
Theorem 1 shows that our SCTD loss achieves an arbitrarily small distillation error when the ODE solver’s step size $\Delta_{t}$ and the segment length $s_{m+1}-s_{m}$ are sufficiently small. By choosing an appropriate segment number $N_{s}$, we obtain a tighter error upper bound, which improves upon the bound $\mathcal{O}(\Delta t)(T)$ in CDS [48] and $\mathcal{O}(\Delta t)(T-e)$ in GCS [22]. This refinement enables more accurate gradients for optimizing the 3D representation $\boldsymbol{\theta}$, leading to higher-quality generative results.
5 Experiments
Figure 4: Qualitative comparison results of two SD-based metheds: DreamFusion [31] and LucidDreamer [23], and two CD-based methods: Consistent3D [48] and ConnectCD [22]. CFG scales are set to 100, 7.5, 20$\sim$40, 7.5, 7.5, respectively. Our approach yields high-quality results within the shortest time. Please zoom in for details.
5.1 Implementation Details
Our algorithm is implemented in PyTorch and trained on an A100 GPU for 5,000 iterations using the Adam optimizer [19]. We utilize Stable Diffusion 2.1 [34] as our foundational generative model and employ Point-E [30] to initialize 3D Gaussians. Hyperparameters related to 3D Gaussian Splatting (3DGS), such as learning rates and camera poses, are set following the guidelines in previous works [23, 22]. We have $t\sim\mathcal{U}(20,500+t_{warm})$, where $t_{warm}$ decreases linearly from 480 to 0 during the first 1,500 epochs.
5.2 Text-to-3D Generation
Qualitative Comparison. We present a collection of 3D assets generated by SegmentDreamer from different camera views in Fig. 1. Our algorithm effectively produces high-quality, text-aligned 3D assets. To highlight the superior generative capability of SegmentDreamer, we conduct qualitative experiments comparing it with state-of-the-art methods based on Consistent Distillation (CD)—namely, Consistent3D [48] and Connect3D [22]—as well as Score Distillation (SD)-based methods, including DreamFusion [31] and LucidDreamer [23], as illustrated in Fig. 4. For a fair comparison, all methods use Stable Diffusion 2.1 as the foundational generative model. As shown in Fig. 4, both DreamFusion and Consistent3D suffer from severe Janus artifacts and fail to generate accurate conditional outputs. LucidDreamer encounters issues with overexposure and ineffective conditional generation. Connect3D performs well in both conditional generation and color rendering; however, its results still contain undesired artifacts, particularly in the tiger and beagle models (see Fig. 4). Moreover, as a 3DGS-based method, it requires over an hour to generate a single 3D asset, limiting its practical usability. In contrast, SegmentDreamer can generate high-quality results within about 32 minutes using CFG [9] and 38 minutes using Perp-Neg [1], proving its efficiency and superiority. More comparison results are shown in our supplementary material.
We also compare our SCTD loss with existing CD-based losses, namely CDS [48] and GCS [22]. For a fair comparison, we use Stable Diffusion 2.1 as the generative model and keep the random seed, guidance scale, and 3DGS-related parameters consistent across all methods. As shown in Fig. 5, CDS struggles to provide effective conditional guidance at a normal CFG scale, as also noted in [22]. Even at a higher CFG scale, its results remain unsatisfactory due to the lack of cross-consistency in its distillation loss. GCS, meanwhile, suffers from overexposure and artifacts due to excessive conditional guidance. While it attempts to address this with Brightness-Equalized Generation (BEG), artifacts persist (see Fig. 5; zoom in for details). In contrast, our SCTD mitigates ineffective conditional generation, overexposure, and artifacts, striking a better balance between self-consistency and cross-consistency.
Figure 5: Visual comparison results of CDS [48], GCS [22], and GCS+BEG [22]. Artifacts are marked with red circles.
Figure 6: Ablation study of the trajectory segmentation strategy with different segment numer $N_{s}$ and time theshold $t_{\tau}$. Prompt: a DSLR photo of the Imperial State Crown of England.
Figure 7: Visual comparison results before and after performing the approximation strategy proposed in Sec. 4.3. “w/ Approx.” and “w/o Approx.” indicate whether the approximation strategy is used or not. The GFG scale is set to 7.5 for all generations.
Quantitative Comparison. We evaluate the text alignment and human preference of each text-to-3D method as follows: we generate 3D models from 40 prompts selected from the DreamFusion Gallery  [^1] https://dreamfusion3d.github.io/gallery.html, and calculate the average CLIP Score and ImageReward (IR) score across 120 camera views for each model. Following Connect3D [22], we compute the FID score between 40 objects (with 120 rendered views for each) and 50,000 images generated by Stable Diffusion 2.1. As shown in the results, our method achieves the best performance in text alignment, human preference, and fidelity, while requiring the shortest generation time, demonstrating its superior overall performance. The 40 prompts used are provided in the supplementary material.
User Study. We conduct a user study involving 40 volunteers to evaluate the generation quality of each algorithm. Each volunteer is asked to review eight groups of five 360-degree rendered videos and to rank them based on three criteria: text alignment (Q1), object realism (Q2), and object details (Q3; e.g., artifacts, brightness, and the Janus problem). To minimize bias, the order of the videos is randomized. As shown in Tab. 2, users consistently prefer our method across all criteria.
6 Ablation Study
Table 2: Quantitative comparison using CLIP score, IR score, training time (min), FID score, and three user study questions given by 40 volunteers. Bold and underline texts denote the best and second-best performance, respectively.
[…]
DreamFusion[31] 28.47 -0.004 140.84 60 4.73 4.87 4.90
LucidDreamer[23] 29.99 0.006 121.80 45 2.93 2.98 2.93
Consistent3D[48] 30.60 0.004 113.61 140 4.14 3.88 3.93
ConnectCD [22] 30.73 0.018 112.61 80 1.63 1.80 1.93
SegmentDreamer 30.88 0.020 110.45 38 1.57 1.47 1.33
[…]
Effect of Segment Number $N_{s}$. We visualize the effect of different segment numbers $N_{s}$ in Fig. 6. As observed, a small $N_{s}$, regardless of the segmentation strategy, consistently results in overly smoothed outputs. As $N_{s}$ increases, the generation quality improves. However, when $N_{s}$ becomes too large (e.g. 10), while more details emerge, the overall model representation becomes vague—a phenomenon also noted in LucidDreamer [23]. We find that $N_{s}=5$ performs well in most cases.
Effect of time threshold $t_{\tau}$. As shown in Fig. 6, a large $t_{\tau}$ leads to overly smoothed results for both segmentation strategies. This is because a limited number of sampling steps in the diffusion process fail to effectively preserve the information of the initial point $\mathbf{z}_{0}$, highlighting the necessity of incorporating multi-step deterministic sampling.
Validation of Approximation Strategy. We analyze the impact of replacing $\boldsymbol{G}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$ with $\mathbf{z}_{s_{m}}$ in Fig. 7. As shown, $\boldsymbol{G}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$ introduces slight overexposure and geometric distortion. In contrast, $\mathbf{z}_{s_{m}}$ not only produces better generation results but also reduces the training time. This suggests a gap between theory and practice, similar to the omission of U-Net Jacobian in DreamFusion [31].
7 Conclusion
We analyze improper conditional guidance in previous Consistency Distillation methods and identify its root cause as an imbalance between self-consistency and cross-consistency in denoising trajectories. Additionally, we find that large distillation error upper bounds limit generative capability. To address this, we propose the Segmented Consistency Trajectory Distillation loss, which improves conditional guidance and tightens the distillation bound. Experiments show that our SegmentDreamer achieves high-quality generation efficiently, advancing distillation-based text-to-3D methods toward practical use.
Limitation. SegmentDreamer mainly focuses on single-instance generation and performs suboptimally in multi-instance generation.
Potential Negative Impact. SegmentDreamer focuses on content synthesis, which may be used to create misleading content.
Acknowledgements
This project is supported by the National Natural Science Foundation of China (12326618, U22A2095), the National Key Research and Development Program of China (2024YFA1011900), the Major Key Project of PCL (PCL2024A06), and the Project of Guangdong Provincial Key Laboratory of Information Security Technology (2023B1212060026).
Supplementary Material
To provide proofs of SegmentDreamer and visual comparisons of state-of-the-arts, this supplementary material includes the following contents:
  • Section A: Guided Consistency Sampling Loss
  • Section B: Flaws in the Consistency Function of Guided Consistency Sampling
  • Figure 1: Oversaturation and Artifacts in GCS
  • Section D: How to Connect SCTD with SDS
  • Section E: Computation Reduction Trick
  • Section F: Proof of Upper Bound of Distillation Error
  • Section G: Visual Comparisons of State-of-the-arts
A Guided Consistency Sampling Loss
The Guided Consistency Sampling (GCS) loss [22] is composed of a compact consistency loss $\mathcal{L}_{\text{CC}}$, a conditional guidance loss $\mathcal{L}_{\text{CG}}$, and a pixel-wise constraint loss $\mathcal{L}_{\text{CP}}$, which are defined by
$\mathcal{L}_{\text{CC}}(\boldsymbol{\theta})=\mathbb{E}[\left\|\boldsymbol{G_{\theta}}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,e,\emptyset)-\boldsymbol{G_{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,e,\emptyset)\right\|_{2}^{2}],$ (1)
$\mathcal{L}_{\text{CG}}(\boldsymbol{\theta})=\mathbb{E}[\left\|\boldsymbol{F_{\theta}}(\mathbf{z}_{e},e,\emptyset)-\boldsymbol{F_{\theta}}(\boldsymbol{G}_{\theta}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,e,\mathbf{y}),e,\emptyset)\right\|_{2}^{2}],$
$\mathcal{L}_{\text{CP}}(\boldsymbol{\theta})\!=\!\mathbb{E}[\left\|\!\boldsymbol{D}(\!F_{\theta}(\mathbf{z}_{e},e,\emptyset))\!\!-\!\!\boldsymbol{D}(\boldsymbol{F_{\theta}}(\boldsymbol{G}_{\theta}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,e,\mathbf{y}),e,\mathbf{y})\!)\!\right\|_{2}^{2}],$
where $e<s<t$, $\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t}$ is estimated by the following trajectory: $\mathbf{z}_{e}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}^{*}\to\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}=\boldsymbol{\Phi}(\mathbf{z}_{e},e,s,\emptyset)\to\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t}=\boldsymbol{\Phi}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{s},s,t,\emptyset)$, $\hat{\mathbf{z}}^{\boldsymbol{\Phi}}_{s}=\boldsymbol{\Phi}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,s,\mathbf{y})$, and $\boldsymbol{D}$ denotes the VAE decoder.
B Flaws in the Consistency Function of Guided Consistency Sampling
As we know, given a well-trained diffusion model $\boldsymbol{\phi}$, there exists an exact solution from timestep $t$ to $e$ [26]:
$\boldsymbol{G}(\mathbf{z}_{t},t,e,\mathbf{y})\!=\!\frac{\alpha_{e}}{\alpha_{t}}\mathbf{z}_{t}+\alpha_{s}\!\int_{\lambda_{t}}^{\lambda_{e}}e^{-\lambda}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t_{\lambda}(\lambda)},t_{\lambda}(\lambda),\mathbf{y})\mathrm{d}\lambda,$ (2)
where $\lambda_{t}=\mathrm{ln}\frac{\alpha_{t}}{\sigma_{t}}$ and $t_{\lambda}$ denotes the inverse function of $\lambda_{t}$. Inspired by [44], we find Eq. 6 suggests that GCS aims to optimize a 3D representation $\boldsymbol{\theta}$ such that $\boldsymbol{G_{\theta}}(\mathbf{z}_{t},t,e,\emptyset)=\boldsymbol{G_{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,e,\emptyset)$ for $\forall t,s,e\in[0,T]$ where $t>s>e$. This implies $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})$, as defined in Eq. 5, is not an approximation but an exact solution learning related to the 3D representation $\boldsymbol{\theta}$, i.e., $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\int_{\lambda_{t}}^{\lambda_{e}}e^{-\lambda}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t_{\lambda}(\lambda)}^{\mathbf{c}},t_{\lambda}(\lambda),\mathbf{y})\mathrm{d}\lambda}{\int_{\lambda_{t}}^{\lambda_{e}}e^{-\lambda}\mathrm{d}\lambda}$. However, dropping the target timestep $e$ in $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})$, as GCS does, is problematic. Suppose we predict both $\mathbf{z}_{e}$ and $\mathbf{z}_{e^{\prime}}$ from $\mathbf{z}_{t}$ where $t>e^{\prime}>e$, we must have
$\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\int_{\lambda_{t}}^{\lambda_{e^{\prime}}}e^{-\lambda}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t_{\lambda}(\lambda)},t_{\lambda}(\lambda),\mathbf{y})\mathrm{d}\lambda}{\int_{\lambda_{t}}^{\lambda_{e^{\prime}}}e^{-\lambda}\mathrm{d}\lambda}$ (3)
$\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\int_{\lambda_{t}}^{\lambda_{e}}e^{-\lambda}\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t_{\lambda}(\lambda)},t_{\lambda}(\lambda),\mathbf{y})\mathrm{d}\lambda}{\int_{\lambda_{t}}^{\lambda_{e}}e^{-\lambda}\mathrm{d}\lambda}.$
[…]
C Oversaturation and Artifacts in GCS
Figure 1: Analysis on self- and cross-consistency in GCS (a), and visual analysis of oversaturation and artifacts (b)-(d).
Fig. 1 shows that the cross-consistency loss $\mathcal{L}_{\text{CG}}+\mathcal{L}_{\text{CP}}$ (with $\mathcal{L}_{\text{CP}}$ dominating) greatly exceeds the self-consistency loss $\mathcal{L}_{\text{CC}}$. As shown in Fig. 1, we scale down $\mathcal{L}_{\text{CP}}$ (c) and observe that the oversaturation (b) is alleviated, but undesirable geometry persists (red circle). This suggests that such oversaturation is brought from an excessively large value of cross-consistency loss, i.e., the “excessive conditional guidance,” which has been analyzed in Sec. 4.1. In contrast, our $\mathcal{L}_{\text{SCTD}}$ (d) successfully addresses the above issues.
D How to Connect SCTD with SDS?
As described in Sec. 4.3, given any subtrajectory $[s_{m},s_{m+1}]$, $\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,s_{m},\mathbf{y}):=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\alpha_{s_{m}}}{\alpha_{t}}\mathbf{z}_{t}-\alpha_{s_{m}}\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda$. Then, we have
$\boldsymbol{\epsilon}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})=\frac{\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\mathbf{z}_{t},t,\mathbf{y})-\frac{\alpha_{s_{m}}}{\alpha_{t}}\mathbf{z}_{t}}{\alpha_{s_{m}}\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda},$ (4)
where $\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$. In this case, $\mathbf{y}$ can represent any prompt embedding, including $\emptyset$. According to Eq. 7, $\mathcal{L}_{\text{SDS}}$ can be further transformed into:
$\hat{\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}=\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})+$ (5)
$\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})+\omega(\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})-\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\emptyset)),$
where $\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\mathbf{z}_{t},t,s,\mathbf{y})$. By substituting Eq. 4 into Eq. 5, we obtain
$\mathcal{L}_{\text{SDS}}(\boldsymbol{\theta})\!=\!\mathbb{E}_{t}[b(t)\!||\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\mathbf{y})$ (6)
$+\omega(\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\emptyset)\!-\!\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,\mathbf{y}))\!-\!\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})$
$+\frac{\alpha_{s_{m}}}{\alpha_{t}}\mathbf{z}_{t}-\alpha_{s_{m}}\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda\boldsymbol{\epsilon}||_{2}^{2}],$
where $b(t)=\frac{\omega(t)}{(\alpha_{s_{m}}\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda)^{2}}$. Furthermore,
$\frac{\alpha_{s_{m}}}{\alpha_{t}}\mathbf{z}_{t}-\alpha_{s_{m}}\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda\boldsymbol{\epsilon}$ (7)
$=\frac{\alpha_{s_{m}}}{\alpha_{t}}(\alpha_{t}\mathbf{z}_{0}^{\mathbf{c}}+\sigma_{t}\boldsymbol{\epsilon})+\alpha_{s_{m}}(e^{\lambda_{s_{m}}-\lambda_{t}}-1)\boldsymbol{\epsilon}$
$=\alpha_{s_{m}}\mathbf{z}_{0}+\sigma_{s_{m}}\boldsymbol{\epsilon}$
$=\mathbf{z}_{s_{m}}$
Substituting Eq. 7 into Eq. 6, we can obtain Eq. 8. Proof is completed.
E Computation Reduction Trick
Eq. 7 can be transformed into
$\hat{\boldsymbol{\epsilon}}_{\boldsymbol{\phi}}(\mathbf{z}_{t},t,\mathbf{y})\!-\!\boldsymbol{\epsilon}\!=\!\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})\!-\!\boldsymbol{\epsilon}\!+\!\omega(\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})\!-\!\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\emptyset))$ (8)
$=\!\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\emptyset)\!-\!\boldsymbol{\epsilon}\!+\!(\omega+1)(\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\mathbf{y})\!-\!\boldsymbol{\epsilon_{\phi}}(\mathbf{z}_{t},t,\emptyset))$
Based on Eq. 8, we can readily derive Eq. 9 by following the procedure described in App. D.
F Proof of Upper Bound of Distillation Error
Before proving Theorem 1, we first give the following lemma:
Lemma 1.
Given a sub-trajectory $[s_{m},s_{m+1}]$, let $\Delta t=max_{t,s\in[s_{m},s_{m+1})}\{|t-s|\}$. We assume $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}$ satisfies the Lipschitz condition and the ODE solver has local error uniformally bounded by $O(t-s)^{p+1}$ with $p\geq 1$. If $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\emptyset)=\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\emptyset)$ for $\forall t,s\in[s_{m},s_{m+1}]$, we have
$\sup_{t,s\in[s_{m},s_{m+1})}\!\!\!\!\!||\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,\mathbf{y})-\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,s_{m},\mathbf{y})||$ (9)
$=O((\Delta t)^{p})(s_{m+1}-s_{m}).$
Proof.
The proof is based on [40, 44, 22, 48]. Let
$\boldsymbol{e}_{n-1}:=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,s_{m},\mathbf{y}),$ (10)
where $\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\mathbf{z}_{s_{m}},s_{m},s,\mathbf{y})$. According to the condition, we have
$\boldsymbol{e}_{n}=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})-\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,s_{m},\mathbf{y})$ (11)
$=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})$
$+\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,s_{m},\mathbf{y})$
$=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})+\boldsymbol{e}_{n-1},$
Provided that $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}$ satisfies $L$-Lipschitz condition, we have
$||e_{n}||=||e_{n-1}+\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})||$ (12)
$\leq||e_{n-1}||+||\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})-\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})||$
$\leq\boldsymbol{e}_{n-1}+L||\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}-\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}||$
$\leq\boldsymbol{e}_{n-1}+L\cdot O((t-s)^{p+1})$
$\leq\boldsymbol{e}_{n-1}+L(t-s)\cdot O((\Delta t)^{p}).$
Besides, according to the boundray condition,
$\boldsymbol{e}_{s_{m}}=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\hat{\mathbf{z}}_{s_{m}}^{\boldsymbol{\Phi}},s_{m},\mathbf{y})-\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{s_{m}}^{\boldsymbol{\Phi}},s_{m},s_{m},\mathbf{y})$ (13)
$=\hat{\mathbf{z}}_{s_{m}}^{\boldsymbol{\Phi}}-\tilde{\mathbf{z}}_{s_{m}}^{\boldsymbol{\Phi}}=\mathbf{z}_{s_{m}}-\mathbf{z}_{s_{m}}=0,$
Therefore,
$||\boldsymbol{e}_{n}||\leq||\boldsymbol{e}_{s_{m}}||+L\sum_{t_{i},t_{i-1}\in[s_{m},s_{m+1}]}(t_{i}-t_{i-1})O((\Delta t)^{p})$ (14)
$=O((\Delta t)^{p})\cdot(s_{m+1}-s_{m}).$
The proof is completed. ∎
According to Eq. 12, one can ideally optimize a 3D model $\boldsymbol{\theta}$ such that $\mathbf{z}_{s_{m}}=\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$. In this case, we have $\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,s_{m},\mathbf{y})=\mathbf{z}_{s_{m}}^{data}$, where $\mathbf{z}_{s_{m}}^{data}=\alpha_{s_{m}}\mathbf{z}^{data}+\sigma_{s_{m}}\boldsymbol{\epsilon}^{*}$. Based on this, we have
$||\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\tilde{\mathbf{z}}^{\boldsymbol{\Phi}}_{t},t,\mathbf{y})-\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,s_{m},\mathbf{y})||\!=\!||\mathbf{z}_{s_{m}}\!-\!\mathbf{z}^{data}_{s_{m}}||\!=\!||\mathbf{z}_{0}\!-\!\mathbf{z}^{data}||.$ (15)
Since we use a first-order ODE solver to implement $\boldsymbol{\Phi}$, we have
$\sup_{t,s\in[s_{m},s_{m+1})}||\mathbf{z}_{0}-\mathbf{z}^{data}||\\ =\mathcal{O}(\Delta t)(s_{m+1}-s_{m}).$ (16)
The proof is completed.
G Visual Comparisons of State-of-the-arts
We also present additional visual comparisons with state-of-the-art methods, as shown in Fig. 2 and Fig. 3. We provide additional qualitative comparisons against DreamFusion [31], LucidDreamer [23], Consistent3D [48], ConnectCD [22], Magic3D [24], Fantasia3D [3], and CSD [53]. As shown, our method clearly outperforms others visually.
Figure 2: Additional qualitative comparisons with DreamFusion [31], LucidDreamer [23], Consistent3D [48], and ConnectCD [22]. CFG scales are set to 100, 7.5, 20$\sim$40, 7.5, 7.5, respectively. Our approach yields results with high quality. Please zoom in for details.
Figure 3: Additional qualitative comparisons with Magic3D [31], Fantasia3D [23], and CSD [48]. Our approach yields results with high quality. Please zoom in for details.
Table 1: 40 prompts for evaluation.
Column 1 Column 2
1. A cat with a mullet 21. A blue motorcycle
2. A pig wearing a backpack 22. Michelangelo style statue of an astronaut
3. A DSLR photo of an origami crane 23. A DSLR photo of a chow chow puppy
4. A photo of a mouse playing the tuba 24. A DSLR photo of cats wearing eyeglasses
5. An orange road bike 25. A red panda
6. A ripe strawberry 26. A DSLR photo of an elephant skull
7. A DSLR photo of the Imperial State Crown of England 27. An amigurumi bulldozer
8. A photo of a wizard raccoon casting a spell 28. A typewriter
9. A DSLR photo of a corgi wearing a top hat 29. A red-eyed tree frog, low poly
10. A rabbit, animated movie character, high-detail 3D model 30. A DSLR photo of a chimpanzee wearing headphones
11. A panda rowing a boat 31. A robot made out of vegetables
12. A highly detailed sand castle 32. A DSLR photo of a red rotary telephone
13. A DSLR photo of a chimpanzee dressed like Henry VIII king of England 33. A DSLR photo of a blue lobster
14. A photo of a skiing penguin wearing a puffy jacket, highly realistic DSLR photo 34. A DSLR photo of a squirrel flying a biplane
15. A blue poison-dart frog sitting on a water lily 35. A DSLR photo of a baby dragon hatching out of a stone egg
16. A DSLR photo of a bear dressed in medieval armor 36. A DSLR photo of a bear dancing ballet
17. A DSLR photo of a squirrel dressed like a clown 37. A plate of delicious tacos
18. A plush toy of a corgi nurse 38. A DSLR photo of a car made out of cheese
19. A humanoid robot playing the violin 39. A yellow school bus
20. A DSLR photo of a bear dressed as a lumberjack 40. A DSLR photo of a shiny beetle